A circular loop in the plane of the paper lies in a 0.90 T magnetic field pointing into the paper. The resistance of the coil is 1.80 Ω. If the loop's diameter changes from 21.5 cm to 9.00 cm in 0.45 s, what is the average induced current in milli-amps? incorrect answers - 13.635 mA and 35.13 mA
Answer #1

here,

magnetic field, b = 0.90 T

resistance, R = 1.80 ohms

change in diameter, d = 21.5 - 9 = 12.5 cm = 0.125 m

time, t = 0.45 s

From faraday law , induced emf is given by expression
induced EMF = rate of change of flux * area
induced EMF = b/t * area
induced EMF = 0.9/0.45 * pi * (0.125/2)^2
induced EMF = 0.025 V

induced Current, I = EMF / R
induced Current, I = 0.025 / 1.8
induced Current, I = 0.014 A or 14 mA